OpenStudy (anonymous):

Adrianne and Peter run a catering company. They have two different pricing options. The first option offers $150 for their services plus $25 per meal they serve. The second option allows the client to only spend $120, but they must pay $40 for each meal served. How many people are served when the total price is the same for both plans?

OpenStudy (anonymous):

lalalala

OpenStudy (anonymous):

i think it's 4 but i'm not sure.

OpenStudy (anonymous):

Let x be the number of meals where the total price is the same for both plans. 150+25x=120+40x 150+25x-120=120+40x-120 30+25x-25x=40x-25x 30=15x x=2